# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

### v3

@@ 3.3. REPS Design Advantages
REPS is simple for cost-efficient hardware support, as it does not require any change of the packet headers format or existing network components. Moreover, its code is short and simple to implement and understand. REPS works best with per-packet ACKs, but we show in Section 4.5.2 that it still performs well even with ACK coalescing.
@@
A key advantage of REPS is that it does not need to track per-EV metrics and statistics. As will be discussed in Section 4.5.1, achieving good performance with OPS requires a relatively large EVS. If OPS were to maintain metrics for each EV, the memory overhead would be excessive for a hardware NIC implementation, e.g., 8 KiB to store 1 byte per entropy value for an EVS with 8K EVs. However, REPS only needs a fixed number of bytes in memory regardless of the EVS size. More specifically, as detailed in Table 1, REPS requires only around $\approx 25$ bytes. Moreover, even when constrained to a small EVS, REPS is still able to perform well (Section 4.5.1), which can further reduce REPS’ memory footprint by 1 byte since Table 1 assumed 16 bits per EV.
There is a subtle observation as to why REPS can achieve a great performance without needing a lot of state: while the REPS buffer is used to cache good entropies, it is really only useful in certain scenarios like when receiving a burst of ACKs or during freezing mode. In reality, most of REPS’ state is on the wire, stored in the inflight data and ACK packets, which will inform REPS about the good paths in the network.
@@
 Number Valid Entropies  (numberValidEVs) 8
@@
Table 1. Per-connection memory footprint of REPS.
@@
The general approach of REPS is that it only keeps track of good paths and avoids keeping statistics on congested or failing paths. This approach enables it to promptly load balance away from a congested link or failing link as, especially for the latter, it is never going to take a random guess once a link is failing. Any alternative method that tries to avoid selecting a failing path by tracking bad EVs would need to keep records of all the EVs that map to that path for a given flow, which would involve tracking not only the failing EVs but also all those still in flight.
@@ 4. Evaluation
Our evaluation consists of simulations that stress test REPS at large scale with a number of workloads. We also evaluate REPS at a meaningful scale on real hardware. Our goal is to answer the following research questions:
  • Does REPS offer an advantage over OPS under baseline conditions in a healthy network with symmetrical topology? (Section 4.3.1 and 4.4.1)
  • Does REPS perform well under network topology asymmetries? (Section 4.3.2 and 4.4.2)
  • Is REPS able to quickly recover from failures? (Section 4.3.3 and 4.4.3)
  • Can REPS work well even with different network parameters and settings? (Section 4.5)
@@ 4.1. Evaluation Setup
We compare REPS with OPS. In the large-scale simulations, we also compare REPS with ECMP (Hopps, 2009), PLB (Qureshi et al., 2022), MPRDMA (Lu et al., 2018), Flowlet Switching (Vanini et al., 2017), MPTCP (Raiciu et al., 2011), a bitmap approach where we keep per EV statistics similarly to STrack (Le et al., 2024), and adaptive RoCE by NVIDIA (NVIDIA, 2024). We configure PLB to have more aggressive parameters similarly to FlowBender to improve its performance (Kabbani et al., 2014). For Flowlet Switching we set an aggressive flowlet timeout at half of the RTT. For MPTCP, we divide each message into 8 subflows and route each one with different EVs similarly to what happens with multiple QPs (Queue Pairs) (Gangidi et al., 2024).
@@
In all simulated baseline runs, we use the same DCTCP (Alizadeh et al., 2010a) variant used in MPRDMA (Lu et al., 2018). It applies per-ACK congestion window updates, allows the receiver to accept and acknowledge out-of-order packets, and reduces the congestion window by one MTU in case of packet drops. In the FPGA-based experimentation, we use a similar but proprietary CC algorithm that relies on ECN marking, congestion notification packets, and per-flow congestion window adjustments.
@@
Regardless of the workloads that we discuss in Section 4.2, in the evaluation we simulate 3 different scenarios: (1) healthy symmetric topology network conditions, (2) asymmetric network conditions (e.g., due to failures, in-order ECMP-hashed background traffic that increase load on specific paths or incremental deployments), and (3) a network encountering various failures. We focus on the most relevant ones for real-world deployments (Gangidi et al., 2024) and we report some of the remaining ones in Appendix D.
@@
We implement REPS by extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and 128 nodes and with different levels of oversubscription ranging from 1:1 (no oversubscription) to 4:1. We test 2- and 3-tier fat trees (TOR or Top-of-rack as T0, Aggregate as T1 and Core T2). Such topologies are commonly deployed in production datacenters designed for distributed training (Singh et al., 2015; Gangidi et al., 2024).
@@
REPS-FPGA::
To demonstrate the effectiveness and resilience of entropy recycling in a real network environment, we also evaluate REPS in an end-to-end setting using a modified production-grade FPGA-based RDMA-capable NIC. Our testbed consists of a two-tier fat-tree Ethernet network with 100G NICs and 12.8T switches. The default MTU for our FPGA NICs is 8KB and typical RTT incorporating NIC buffer delay and ACK processing through T0 and T1 are in the order of 10 and 15 us, respectively.
@@ 4.2. Workloads
We evaluate REPS on a mix of synthetic benchmarks, real datacenter traces, and distributed training collectives.
@@
consists of (1) incast, (2) permutation, and (3) tornado traffic patterns. Incast happens when multiple senders simultaneously send to one receiver. It is very common in storage workloads (Zhang et al., 2011; Chen et al., 2009) but also, with a small incast degree, in distributed training (Gangidi et al., 2024). In the permutation pattern, each node sends to a random receiver, and we ensure that each node is sending and receiving to exactly one node (Alizadeh et al., 2010a). The tornado pattern is a special case of the permutation where each node sends to its ”twin” node in the other half of the tree. For example, with 128 nodes, node 0 would send to 64 and vice-versa, node 1 to 64 and so on. Tornado is an important worst case for load balancing, as each packet is required to traverse the full tree (Prisacari et al., 2013).
@@
We use real datacenter traces from similar previous work (Alizadeh et al., 2010a; Zhang et al., 2017). We use a series of traces used for web search in production clusters. In such distribution the majority of flows are quite small (less than 100 KB) while a small number of flows are large. For each node we select randomly the receiver and run the simulation for 5 ms. More details are available in Appendix E.
@@
We show simulated results for two commonly used collectives in AI training: the AllReduce implemented via the ring and butterfly algorithm (Li et al., 2020), the AllToAll implemented using an algorithm where we limit the number of parallel connections per node ($n$ connections) (Hoefler et al., 2022; Naumov et al., 2019).
Our baseline traffic for REPS-FPGA consists of 128 endpoints under two T0 switches continuously performing 4 MB ring-based AllReduce collective operations, with the logical ring laid out such that all connections traverse the T1 spine to maximize the pressure on the spine of the topology.
@@ 4.3. Simulation Results
We conduct a detailed analysis of REPS behavior for each network condition (see Sec. 4). We first examine a specific case in depth and then summarize key takeaways.
@@ 4.3.1. Healthy Symmetric Network Conditions
In this section we evaluate the performance of REPS in a simple setting where there is no oversubscription and there are no failures, meaning the network is perfectly symmetrical. Intuitively, this seems the best situation for oblivious packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we will see later in this section and based on our simple theoretical model in Section 5, this is not the case as REPS still offers an up to 25% advantage over OPS. This is because of ECMP collisions that still happen with OPS. While over long period of time, each link will be evenly used, there will still be short-term collisions happening that will increase and decrease the link utilization of certain links.
Figure 1. A tornado workload with 16 MiB messages with CC disabled and using OPS and REPS as load balancers.
Figure 2. REPS performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives.
Figure 3. REPS vs. OPS in a 32 MiB message send.
@@
In the case of OPS, we can see that due to the random nature of it, queues are created over time and that the link utilization of each port, at small timeframes, sometimes goes significantly above and below it (15% more or less). This shows that while OPS does still a decent job at completing the workload close to the ideal completion time, it does inevitably create unpredictable queues (potentially even exceeding $K_{max}$ and causing drops) over short time periods. Such queues will cause the CC to kick in and slightly reduce the sending rate and average output port utilization.
In the same configuration with REPS, we notice a major difference in the bottom plot of Figure 1. In particular REPS converge quickly to a configuration where each queue is kept below $K_{min}$ (note that the only guarantee here is that all queues will be below $K_{min}$, not necessarily all at the same value). At the same time, we can also notice that all the ports converge to the perfect selection rate of 400 Gbps. While the overall completion time is only about 4% better than OPS, the smaller queues provide a better guarantees for system low latency traffic. Moreover, as we can see in Figure 2, this gap expands as we increase the message size.
Looking at the port selection rate over the entire run of the workload for OPS vs. REPS, we observe that they are nearly equivalent. This is again because the problem is with the short-term collisions that OPS can experience at microscopic scale. Finally, we note that, while the main advantages of REPS are not experienced with these perfectly symmetric scenarios, we consistently observe lower max out-of-order distance thanks to its improved stability.
@@
We now focus more on the overall view comparing REPS with all the other state of the art algorithms in a series of benchmarks.
In Figure 2, we visualize a summary of the performance of the various algorithms by looking at the runtime of the workloads (max FCT). As expected, in the case of incast, the performance is driven almost exclusively by the CC and, hence, we do not see any major difference between all the load balancers and even ECMP performs well. However, once we move to permutation and tornado workloads, ECMP collisions start to drastically reduce the performance of ECMP. In most cases REPS outperforms all the other algorithms.
In the tornado case, Adaptive RoCE is able to match REPS since this is the ideal scenario for it: REPS, unlike Adaptive RoCE, still needs to guess during its initial BDP worth of packets. On the other hand, REPS outperforms Adaptive RoCE in the permutation pattern where taking a local best decision might not always lead to the best global outcome. We also see the difference between algorithms that were designed to reduce the number of out-of-order packets versus algorithms that do not have such hard constraints. Additionally, there is a distinction between algorithms that operate at packet-level granularity, such as REPS, OPS, BitMap, and MPRDMA, and those that operate at a coarser granularity, such as Flowlet and PLB.
Figure 4. REPS performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives and an asymmetric network due to 2% of the TOR uplinks being offline.
For datacenter traces, we analyze the results for different load levels, ranging from 40% to 100%. Here we observe again, the clear difference between per-packet algorithms and less granular options. Even at higher load REPS is able to work well with a 5% advantage over OPS.
We also run distributed collectives and report their completion times. We observe how, by design, the ring AllReduce has the same performance regardless of the load balancing algorithm utilized. This is because, due to its ring design, there is no opportunity for congestion to accumulate. In AllToAll, REPS gets an up to 20% advantage over the alternatives.
@@ 4.3.2. Asymmetric Network Conditions
We evaluate REPS under different scenarios where some degree of asymmetry is created. We focus on two scenarios: 1) the network is not perfectly symmetrical because of some missing (or degraded) cables, 2) there is some background traffic in the network that is using ECMP routing.
@@
We visualize this problem with a simple scenario where we have a switch with $n$ input and output ports and $n$ flows active, each from a different source sending a 32 MiB message. To create an asymmetry, we reduce one of the uplinks speed to 200 Gbps while all the other links remain at 400 Gbps. In Figure 3 we visualize the output port utilization rate for OPS and REPS.
We observe that while OPS chooses each port equally, irrespective of its actual bandwidth, REPS eventually converges to a stable configuration where the slower uplink is used less frequently. This results in both stabler queues but, more importantly, a much faster completion time ($1400\;\mu\text{s}$ for OPS and $756\;\mu\text{s}$ for REPS).
Figure 5. Synthetic traffic with background ECMP traffic.
@@
We now shift our focus to more general results when encountering asymmetries in a network. For space constraints, we focus mostly on the case where some of the links have a lower sending rate. In our first experiment we run synthetic benchmarks where $2\%$ of the TOR uplinks, chosen randomly, have been downgraded to 200 Gbps. In Figure 4, we can see results similar to before where REPS gets an up to 350% advantage over ECMP and 10% advantage over the second best algorithm (usually BitMap). In the DC traces we can see a higher difference due to the asymmetry in the network. At 100% load REPS gets a 25% advantage over the second best algorithm and a 1000% advantage over ECMP. We show the results for several AI collectives. We note how for AllToAll REPS keep a small but significant advantage and in the AllReduce a sizable 50% advantage over the second best performing algorithm.
Figure 6. REPS performance under different failure modes in a 8 MiB permutation, DC traces at 100% load and a ring AllReduce.
In Figure 5, we showcase one example of REPS sharing traffic together with background ECMP traffic (we assume 10% of the traffic is ECMP). In this case, REPS: 1) shifts REPS traffic away from ECMP traffic in order to not slow down REPS traffic, 2) helps background traffic by ensuring that it will not be slowed down by REPS traffic. This also highlights the possibility of incrementally deploying REPS on ECMP-base systems.
Finally, we note that WCMP (Zhou et al., 2014) could be used to enhance ECMP performance in the case of a topology with known asymmetries, but would not help as much in the case of unpredictable mixed traffic or sudden temporary asymmetries.
Figure 7. REPS vs. OPS in a 32 MiB permutation with two cables’ failure (a shorter one and a longer one).

### v4

@@ 3.3. ARCANE Design Advantages
The simplicity of ARCANE enables cost-efficient hardware support, as it does not require any change of the packet headers format or existing network components. Moreover, its code is short and easy to implement and understand. ARCANE works best with per-packet ACKs, but we show in Section 4.5.1 that it still performs well even with ACK coalescing.
@@
A key advantage of ARCANE is that it does not need to track per-EV metrics and statistics. As will be discussed in Section 4.5.2, achieving good performance with OPS requires a relatively large EVS. If OPS were to maintain metrics for each EV, the memory overhead would be excessive for a hardware NIC implementation, e.g., 64 Kib to store 1 bit per entropy value for an EVS with 64K EVs. However, ARCANE only needs a fixed number of bytes in memory. More specifically, as detailed in Table 1, ARCANE requires only around $\approx 25$ bytes. Moreover, even when constrained to a small EVS, ARCANE is still able to perform well (Section 4.5.2), which can further reduce ARCANE’ memory footprint by 1 byte since Table 1 assumed 16 bits per EV.
There is a subtle observation as to why ARCANE can achieve a great performance without needing a lot of state: while the ARCANE buffer is used to cache good entropies, it is really only useful in certain scenarios like when receiving a burst of ACKs or during freezing mode. In reality, most of ARCANE’ state is on the wire, stored in the inflight data and ACK packets, which will inform ARCANE about the good paths in the network.
@@
 Number Valid Entropies  (numberOfValidEVs) 8
@@
Table 1. Per-connection memory footprint of ARCANE.
@@
The general approach of ARCANE is that it only keeps track of good paths and avoids keeping statistics on congested or failing paths. This approach enables it to promptly load balance away from a congested link or failing link as, especially for the latter, it is never going to take a random guess once a link is failing. Any alternative method that tries to avoid selecting a failing path by tracking bad EVs would need to keep records of all the EVs that map to that path for a given flow, which would involve tracking not only the failing EVs but also all those still in flight.
@@ 4. Evaluation
Our evaluation includes both large-scale simulations designed to stress test ARCANE under a variety of workloads, and experiments on real hardware at meaningful scale. The goal is to answer the following research questions:
  • Does ARCANE outperform state-of-the-art load balancers in a healthy network with a symmetric topology? (Sections 4.3.1 and 4.4.1)
  • How does ARCANE perform in the presence of topology asymmetries or when coexisting with non-ARCANE traffic? (Sections 4.3.2 and 4.4.2)
  • Can ARCANE recover quickly from failures? Is it resilient to severe or widespread failures? (Sections 4.3.3 and 4.4.3)
  • Does ARCANE maintain performance under varying network parameters and configurations? (Section 4.5)
@@ 4.1. Evaluation Setup
In large-scale simulations, we compare ARCANE with ECMP (Hopps, 2009), OPS (Dixit et al., 2013), PLB (Qureshi et al., 2022), MPRDMA (Lu et al., 2018), Flowlet Switching (Vanini et al., 2017), MPTCP (Raiciu et al., 2011), a bitmap approach where we keep per EV statistics similarly to STrack (Le et al., 2024), and adaptive RoCE by NVIDIA (NVIDIA, 2024). We note that this list includes both sender-based approaches and in-network switch-based approaches. We configure PLB to have more aggressive parameters similarly to FlowBender to improve its performance (Kabbani et al., 2014). For Flowlet Switching we set an aggressive flowlet timeout at half of the RTT. For MPTCP, we divide each message into 8 subflows and route each one individually similarly to using multiple QPs (Queue Pairs) (Gangidi et al., 2024).
@@
In all simulated baseline runs, we use the same DCTCP (Alizadeh et al., 2010) variant used in MPRDMA (Lu et al., 2018). It applies per-ACK congestion window updates, allows the receiver to accept and acknowledge out-of-order packets, and reduces the congestion window by one MTU in case of packet drops. In the FPGA-based experimentation, we use a similar but proprietary CC algorithm that relies on ECN marking, congestion notification packets, and per-flow congestion window adjustments.
@@
Regardless of the workloads that we discuss in Section 4.2, in the evaluation we simulate 3 different scenarios: (1) healthy symmetric network conditions, (2) asymmetric network conditions, which can be due to asymmetric path bandwidth or due to ECMP-hashed background traffic that increase load on specific paths (a scenario that can happen in incremental deployments of ARCANE), and (3) a network encountering various failures modes. We focus on the most relevant failure scenarios for real-world deployments (Gangidi et al., 2024) and we report the remaining ones in Appendix D.
@@
We implement ARCANE by extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and 128 nodes and with different levels of oversubscription ranging from 1:1 (no oversubscription) to 4:1. We test 2- and 3-tier fat trees (TOR or Top-of-rack as T0, Aggregate as T1 and Core T2). Such topologies are commonly deployed in production datacenters designed for distributed training (Singh et al., 2015; Gangidi et al., 2024).
@@
ARCANE-FPGA::
To demonstrate the effectiveness and resilience of entropy recycling in a real network environment, we also evaluate ARCANE in an end-to-end setting. We use a modified production-grade FPGA-based RDMA-capable NIC running a custom transport protocol capable of handling OOO packets by directly placing the payload in memory via RDMA. The transport implementation relies on SACKs that carry bitmaps (256-bit-wide in these experiments) relaying packet delivery updates to track completions and handle selective retransmissions. Each connection maintains an 8-entry-deep ARCANE buffer that stores UDP source ports (2B). Our experimental setup is configured to support up to 256 connections for a total memory footprint of 4KB. Given that only a single connection is active at a time, all ARCANE buffers can be hosted in a single SRAM memory and the associated logic to manage ARCANE buffers can be multiplexed across all connections, occupying negligible resources (¡0.04% of the total device logic resources).
The testbed consists of a two-tier fat-tree Ethernet network with 100G NICs and 12.8T switches. The default MTU for our FPGA NICs is 8KB and typical RTT incorporating NIC buffer delay and ACK processing through T0 and T1 are in the order of 10 and 15 us, respectively.
@@ 4.2. Workloads
We evaluate ARCANE on a mix of synthetic benchmarks, real datacenter traces, and distributed training collectives.
@@
consists of (1) incast, (2) permutation, and (3) tornado traffic patterns. Incast happens when multiple senders simultaneously send to one receiver. It is very common in storage workloads (Zhang et al., 2011; Chen et al., 2009) but also, with a small incast degree, in distributed training (Gangidi et al., 2024). In the permutation pattern, each node sends to a random receiver, and we ensure that each node is sending and receiving to exactly one node (Alizadeh et al., 2010). The tornado pattern is a special case of the permutation where each node sends to its ”twin” node in the other half of the tree. For example, with 128 nodes, node 0 would send to 64 and vice-versa, node 1 to 64 and so on. Tornado is an important worst case for load balancing, as each packet is required to traverse the full tree (Prisacari et al., 2013).
@@
We use real datacenter traces from similar previous work (Alizadeh et al., 2010; Zhang et al., 2017). We use a series of traces used for web search in production clusters. In such distribution the majority of flows are quite small (less than 100 KB) while a small number of flows are large. For each node we select randomly the receiver and run the simulation for 5 ms. More details are available in Appendix E.
@@
We show simulated results for two commonly used collectives in AI training: the Allreduce implemented via the ring and butterfly algorithm (Li et al., 2020), the Alltoall implemented using an algorithm where we limit the number of parallel connections per node ($n$ connections) (Hoefler et al., 2022a; Naumov et al., 2019).
Our baseline traffic for ARCANE-FPGA consists of 128 endpoints under two T0 switches continuously performing 4 MB ring-based AllReduce collective operations, with the logical ring laid out such that all connections traverse the T1 spine to maximize the pressure on the spine of the topology.
@@ 4.3. Simulation Results
We conduct a detailed analysis of ARCANE behavior for each network condition (see Section 4). We first examine a specific case in depth and then summarize key takeaways.
@@ 4.3.1. Healthy Symmetric Network Conditions
In this Section we evaluate the performance of ARCANE in a simple setting where there is no oversubscription and there are no failures, meaning the network is perfectly symmetrical. Intuitively, this seems the best situation for oblivious packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we will see later in this section and based on our theoretical model in Section 5, this is not the case as ARCANE still offers an up to 25% advantage over OPS. This is because of ECMP collisions that still happen with OPS. While over long period of time, each link will be evenly used, there will still be short-term collisions happening that will increase and decrease the link utilization of certain links.
Figure 1. A tornado workload with 16 MiB messages using OPS (top) and ARCANE (bottom) as load balancers.
Figure 2. ARCANE performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives.
Figure 3. ARCANE vs. OPS in a 32 MiB message send with an asymmetric topology.
@@
In the case of OPS, we can see that due to the random nature of it, queues are created over time and that the link utilization of each port, at small timeframes, sometimes goes significantly above and below the optimal point (15% more or less). This shows that while OPS does still a decent job at completing the workload close to the ideal completion time, it does inevitably create unpredictable queues (potentially even exceeding $K_{max}$ and causing drops) over short time periods. Such queues will cause the CC to kick in and slightly reduce the sending rate and average output port utilization.
In the same configuration with ARCANE, we notice a major difference in the bottom plot of Figure 1. In particular ARCANE converge quickly to a configuration where each queue is kept below $K_{min}$ (note that the only guarantee here is that all queues will be below $K_{min}$, not necessarily all at the same value). At the same time, we can also notice that all the ports converge to the perfect selection rate of 400 Gbps. While the overall completion time is only about 4% better than OPS, the smaller queues provide a better guarantees for system low latency traffic. Moreover, as we can see in Figure 2, this gap expands as we increase the message size.
Looking at the port selection rate over the entire run of the workload for OPS vs. ARCANE, we observe that they are nearly equivalent. This is, again, because the problem is with the short-term collisions that OPS can experience at microscopic scale.
@@
We now focus more on the overall view comparing ARCANE with all the other state of the art algorithms in a series of benchmarks.
In Figure 2, we visualize a summary of the performance of the various algorithms by looking at the runtime of the workloads (max FCT). As expected, in the case of incast, the performance is driven almost exclusively by the CC and, hence, we do not see any major difference between all the load balancers and even ECMP performs well. However, once we move to permutation and tornado workloads, ECMP collisions start to drastically reduce the performance of ECMP. In most cases ARCANE outperforms all the other algorithms.
In the tornado case, Adaptive RoCE is able to match ARCANE since this is the ideal scenario for it: ARCANE, unlike Adaptive RoCE, still needs to guess during its initial BDP worth of packets. On the other hand, ARCANE outperforms Adaptive RoCE in the permutation pattern where taking a local optimal decision might not always lead to the best global outcome. We also see the difference between algorithms that were designed to reduce the number of out-of-order packets versus algorithms that do not have such hard constraints. Additionally, there is a distinction between algorithms that operate at packet-level granularity, such as ARCANE, OPS, BitMap, and MPRDMA, and those that operate at a coarser granularity, such as Flowlet and PLB.
For datacenter traces, we analyze the results for different load levels, ranging from 40% to 100%. Here we observe again, the clear difference between per-packet algorithms and less granular options. Even at higher load ARCANE is able to work well with a 5% advantage over OPS.
We also run distributed collectives and report their completion times. We observe how, by design, the ring AllReduce has the same performance for most of the load balancing algorithms utilized. This is because, due to its ring design, there is no opportunity for congestion to accumulate. In AllToAll, ARCANE gets an up to 20% advantage over the alternatives.
Figure 4. ARCANE performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives and an asymmetric network due to 2% of the TOR uplinks being offline.
Figure 5. Synthetic traffic with ARCANE and non-ARCANE (ECMP) traffic, emulating a mixed traffic scenario.
@@ 4.3.2. Asymmetric Network Conditions
We evaluate ARCANE under different scenarios where some degree of asymmetry is created. We focus on two scenarios: 1) the network is not perfectly symmetrical because of some missing (or degraded) cables, 2) there is some background traffic in the network that is using ECMP routing, analyzing the coexistence between ARCANE and non-ARCANE traffic.
@@
We visualize this problem with a simple scenario where we have a switch with $n$ input and output ports and $n$ flows active, each from a different source sending a 32 MiB message. To create an asymmetry, we reduce one of the uplinks speed to 200 Gbps while all the other links remain at 400 Gbps. In Figure 3 we visualize the output port utilization rate for OPS and ARCANE.
We observe that while OPS chooses each port equally, irrespective of its actual bandwidth, ARCANE eventually converges to a stable configuration where the slower uplink is used less frequently. This results in both stabler queues but, more importantly, a much faster completion time ($1400\;\mu\text{s}$ for OPS and $756\;\mu\text{s}$ for ARCANE).
@@
We now shift our focus to more general results when encountering asymmetries in a network. For space constraints, we focus mostly on the case where some of the links have a lower sending rate. In our first experiment we run synthetic benchmarks where $2\%$ of the TOR uplinks, chosen randomly, have been downgraded to 200 Gbps. In Figure 4, we can see results similar to before where ARCANE gets an up to 350% advantage over ECMP and 10% advantage over the second best algorithm (usually BitMap). In the DC traces we can see a higher difference due to the asymmetry in the network. At 100% load ARCANE gets a 25% advantage over the second best algorithm and a 1000% advantage over ECMP. We show the results for several AI collectives. We note how for AllToAll ARCANE keep a small but significant advantage and in the AllReduce a sizable 50% advantage over the second best performing algorithm.
In Figure 5, we showcase one example of ARCANE sharing traffic together with background ECMP traffic (we assume 10% of the traffic is ECMP). In this case, ARCANE: 1) shifts ARCANE traffic away from ECMP traffic in order to not slow down ARCANE traffic, 2) helps background traffic by ensuring that it will not be slowed down by ARCANE traffic. This also highlights the possibility of incrementally deploying ARCANE on ECMP-base systems. Finally, we note that WCMP (Zhou et al., 2014) could be used to enhance ECMP performance in the case of a topology with known asymmetries, but would not help as much in the case of unpredictable mixed traffic or sudden temporary asymmetries.
Figure 6. ARCANE vs. OPS in a 32 MiB permutation with two cables’ failure (a shorter one and a longer one).
Figure 7. ARCANE performance under different failure modes in a 8 MiB permutation, DC traces at 100% load and a ring AllReduce.